Missy collects $2 bills and the sum of a number of bills she has and 14 is equal to two times the number of bills plus four. How many $2 bills does Missy have?
Answer by Cromlix(4381)

Number of bills = x
x + 14 + 2x + 4
x - 2x = 4 - 14
-x = -10
x = 10
Missy has 10 $2 bills.
Hope this helps.
:-)